OpenStudy (anonymous):

Simplify: 3x^4 + 2x^4 5x^4 6x^4 5x^12 6x^8

OpenStudy (anonymous):

\[3\spadesuit+2\spadesuit=5\spadesuit\]

OpenStudy (anonymous):

Its 3*x*x*x*x+2*x*x*x*x= 5*x*x*x*x = 5x^4 you just add them!

OpenStudy (anonymous):

Thanks @ocelot40